# Ameliorating transient noise bursts in gravitational-wave searches for intermediate-mass black holes

### v1

@@ II.4 Triggers
[…]
Afterwards, each trigger is ranked Cannon et al. (2015); Hanna et al. (2020b); Tsukada et al. (2023) according to its probability of originating from an actual signal. The goal of the pipeline is then to obtain a list of triggers, ordered by their likelihood $\Lambda$ to be of astrophysical origin.
@@ III.1 Injections and glitches
Here, we describe the simulated IMBH waveforms, and the six different glitch populations from the data set of Gravity Spy during O3 in H1, L1 and V1 Zevin et al. (2017); Glanzer et al. (2023). As Gravity Spy is a classification algorithm, it returns a classification for each glitch. Glitches have been selected with a Gravity Spy confidence level greater than $90\%$, indicating that the model is highly confident that the glitches belong to the assigned class.
@@
Figure 3: Q-transform of GW190521 and example of different glitch classes selected from Gravity Spy at $>90\%$ confidence in LIGO Livingston (L1) Zevin et al. (2017).
  • Injections: For this aim we use IMRPhenomD approximant Khan et al. (2016) with masses $m_{1}\in[50,400]\,\text{M}_{\odot}$ and $m_{2}\in[10,250]\,\text{M}_{\odot}$, and dimensionless spins $\chi_{1},\chi_{2}\in[0,0.99]$, uniformly sampled within the specified range with aligned spins. The distances are uniformly sampled from the range $d\in[10,250]\,$Mpc. The inclination is sampled from a uniform distribution, and the sky-localization is randomized.
  • Blip: these glitches have a characteristic morphology of a symmetric “teardrop” shape in time-frequency in the range $[30,250]\,$Hz with short-durations, $\sim 0.04\,$s Cabero et al. (2019). Due to their abundance and form, these glitches hinder both the unmodeled burst and modelled CBC can searches with particular emphasis on compact binaries with large total mass due to the short duration of their waveforms, among others Abbott et al. (2018b, 2016). Moreover, as their source is still unknown, they cannot be removed from astrophysical searches.
  • Fast__\__Scattering: they appear as short-duration arches ($\sim 0.2-0.3\,$s) in the frequency range $[20-60]\,$Hz. These glitches are strongly correlated with ground motion in range $[0.1-0.3]\,$Hz and $[1-6]\,$Hz, which in turn is associated with thunderstorms and human activity near the detector. Note that this class was added during O3, and it is more abundant in L1 than H1, but not present in Virgo, due to differences in ground motion and detector sensitivity Soni et al. (2021).
  • Koi__\__Fish: their naming comes from their imaginative frontal resemblance to koi fishes. They are also similar to Blip but typically feature higher-SNR than the latter, spanning the frequency range of $\sim 20-1000\,$Hz.
  • Low__\__Frequency__\__Burst: they are short-duration glitches ($\sim 0.25\,$s) in the frequency range $[10-20]\,$ Hz with a distinctive hump shape at the bottom of spectrgrams. These occurrences were prevalent in L1 data during O1 and H1 data in O3a.
@@
  • Whistle: these glitches have a characteristic V, U or W shape at higher frequencies ($\gtrsim$ 128 Hz) with typical durations $\sim 0.25\,$s. They are caused when radio-frequency signals beat with the voltage-controlled oscillators Nuttall et al. (2015).
@@ III.2 Clustering
    Unknown data set: Usually clusters have a centroid with the highest $\rho$, and a set of neighbours with smaller $\rho$. For illustration, in the top panel of Fig. 4 we can observe the SNR of a cluster with this behavior. In a realistic setting, we do not know the GPS time of the centroid of the cluster $t_{cluster}$ a priori, so we define them following a procedure similar to GstLAL. We order all the triggers according to their GPS time. Afterwards, we divide the whole observing run in windows of ±$\pm 1\,$s, starting from the beginning of the run. Within each window, we select the trigger with maximum $\rho$ as the centroid, so that $t_{cluster}=t_{\max{\rho}}$. As before, we select $t_{w}=\pm\,\Delta t$ around $t_{cluster}$ to define the cluster.
@@ III.4 Time shift and false alarm rate
~$\widetilde{\mathcal{F}}(\Lambda^{*})=\frac{\sum_{l=i}^{L}\widehat{N}_{l}(\Lambda\geq\Lambda^{*})}{T_{b}},$ (7)
@@
~$\begin{split}&\widetilde{\mathcal{F}}(\Lambda^{*})=\frac{\sum_{l=i}^{L}\widehat{N}_{l}(\Lambda\geq\Lambda^{*})}{T_{b}T_{f}/T_{f}}\\ &=\frac{\widehat{N}(\Lambda\geq\Lambda^{*})}{T_{f}\widetilde{N}}=\frac{\overline{N}(\Lambda\geq\Lambda^{*})}{T_{f}}.\end{split}$ (8)
@@
Since the offsets of the time shifts should be large enough that each slide can be thought of as an independent experiment, we chose to slide L1 in steps of 333\,3s, and V1 in steps of 666\,6s with respect to H1, which is fixed. We performed enough time shifts to produce $1,000\,$yr of data in triple detector coincidence. Afterwards, the time shifts are binned according to their coincidence type and their time of background.
To minimize the number of false positives, the types of detector time coincidences analyzed in this work are either double or triple. Using Eq. 8, it is necessary to properly measure the time of the search $T_{s}$ and the time of the background $T_{b}$. If we are measuring the time in double-time coincidence, for example, between H1 and L1, it would imply that V1 was down at that time (no V1). Otherwise, if all detectors record data simultaneously, it is considered a triple-time coincidence. In Table 1 we can see an overview of the time of the search $T_{s}$. Since H1L1 (no V1) is extremely small, it is computationally intensive to generate 100010001000\,1000yr of time shifts. Thus, in the interest of time and computational resources we focus on triple-coincident time as we expect a better performance.
@@ IV Methodology
The challenges in GW research require innovative solutions, and ML has emerged as a crucial tool for addressing them due to its adaptability and transversality. The main advantage of ML techniques is their rapidity during the deployment phase, since most of the computations are made during the training stage. In the past few years, researchers have explored different ML applications to GW data analysis, such as detection of CBC George and Huerta (2018); Yan et al. (2022a); Ruder (2016); Gabbard et al. (2018); Sharma et al. (2022); Krastev. (2020); Menéndez-Vázquez et al. (2021); Baltus et al. (2021, 2022); ”Martins and others” (2024); Nousi et al. (2023); Koloniari et al. (2024), burst identification Cavaglia et al. (2020); Antelis et al. (2022); Iess et al. (2020, 2023); López Portilla et al. (2021); López et al. (2021); Meijer et al. (2023); Boudart and Fays (2022); Boudart (2023); Skliris et al. (2024), glitch characterization Biswas et al. (2013); Zevin et al. (2017); Glanzer et al. (2023); Bahaadini et al. (2018); Laguarta et al. (2023) and synthetic data generation Lopez et al. (2022b, c); Powell et al. (2023); Yan et al. (2022b); Dooney et al. (2022, 2024), among others. We refer the interested reader to Cuoco et al. (2021) and Schäfer et al. (2023) for a review.
In Yan et al. (2022a) the authors seek to enhance current state-of-the-art search algorithms with ML applications. Following a similar line of reasoning, and as previously mentioned, we use the triggers generated by the IMBH search of GstLAL during O3 to train a classification algorithm for multiple classes. In this way, we output a statistic to differentiate between IMBH signals and glitches, providing direct information about the nature of these populations.
@@ IV.2 Tackling class imbalance
While we can simulate a large number of GW signals, glitches have a finite nature. Despite their occurrence rate being approximately $\sim 1\,\text{min}^{-1}$, some classes of glitches are more common than others Abbott et al. (2021b). For example, in L1 the LIGO collaboration identified $>10^{4}$ Tomtes, but only $\sim 10^{3}$ Whistles. In Table 2 we show the original size of our training and validation set, which is highly imbalanced. This is problematic for classification ML algorithms, as heavily imbalanced datasets tend to be biased towards the majority class to minimize their loss function. This may wrongly be interpreted as a “good” performance, while in reality, the model is only learning a single class: the largest one. Similarly to Powell et al. (2023), to circumvent this issue we use undersampling or oversampling techniques:
@@
    Oversampling: To oversample we use bootstrapping with replacement, which stochastically resamples the existing data set allowing the same example to appear more than once James et al. (2013). Using this method, we oversample all glitch classes to equal the size of the largest glitch class.
@@
To avoid overfitting we use K-fold cross-validation Bengio and Grandvalet (2004). This method partitions the data set in $K$ subsets or folds (see Fig. 6). The model is evaluated $K$ times, using a different fold as validation in each iteration, while training on the rest of the folds. This allows for a more robust estimate of the model’s performance, as it accounts for the variability in the data and limits the potential for overfitting. In particular, we chose $K=9$ as it is a trade-off between computational complexity and performance.
@@ IV.3 Training model
The unit of an MLP is the perceptron, a mathematical simplification of human neurons Rosenblatt (1958). In this unit, the information is given by a $n$-vector $x=[x_{1},\dots,x_{n}]$, that travels in the neuron and interacts multiplicatively with the weights, $w=[w_{1},\dots,w_{n}]$. The goal of the perceptron is to learn the value of the weights, which controls the strength of the influence of the previous neuron. In this simplistic model, the information is summed, such that if the sum is above a certain threshold, the neuron creates or fires a signal. This firing rate is modelled as the activation function $f:\mathbb{R}\rightarrow\mathbb{R}$, which is non-linear. Thus, we can mathematically define the $a$ output, known as activation as,
@@
We implemented the MLP structure in PyTorch Paszke et al. (2019). After several experiments, we selected the best-performing architecture from our results: it consists of 3 hidden layers of 350 each, using the ReLU activation function Agarap (2018), except the output layer that uses Softmax activation function, suitable for multi-class classification, since it is a multi-class task. This MLP architecture is common for each GW detector, but they have been trained separately, i.e. it is a single-detector classifier. To optimize the networks we use cross-entropy loss function, and Adam optimizer Kingma and Ba (2014). Moreover, to adjust the number of epochs and avoid overfitting we implemented an early stopping algorithm Yao et al. (2007). We define an epoch as the number of times the network has passed through the whole training and validation data set. Thus, the early stopping algorithm calculates the difference in validation accuracy $\mathcal{A}_{val}$ between the current epoch, $e$, and the best epoch, $e^{\text{best}}$. The training process finishes if
@@
during 150 epochs, with $\epsilon=0.0001$. For the learning rate, we use an adaptive learning rate built-in the PyTorch function ReduceLROnPlateau. Setting the initial learning rate to $10^{-3}$, if the validation accuracy remains constant after 100100100\,100epochs, the learning rate decreases $10\%$. The combination of these methods proved to increase the performance while decreasing the time needed for fine-tuning.
@@ IV.4 Time coincident tracks
Even though the MLP receives direct information from the Injection class and various glitch classes, misclassifications are still common due to the simplicity of the feature vector (see Section IV.1). Nonetheless, if a potential GW signal is detected in multiple detectors, the likelihood of it being of astronomical origin increases significantly. Hence, it is standard in GW searches to evaluate triggers that happened within the light time travel between detectors: 101010\,10ms between H1 and L1, 272727\,27ms between H1 and V1, and 262626\,26ms between L1 and V1. An additional 555\,5ms is considered for statistical fluctuations Sachdev et al. (2019).
@@
In this work, the approach is inherently different, as we do not deal with individual triggers but with averaged tracks. Thus, two tracks from different detectors are considered time coincident if the average time of the triggers within each track falls within the light travel time and time fluctuations. Nevertheless, even if two tracks are time coincident, there is no guarantee that they originate from the same signal. Therefore, two tracks are only considered time coincident if they have at least one common trigger. If tracks from different detectors coincide in time, we assign to the coincidence a probability equal to the harmonic mean, one of Pythagorean means Larry and Niven (1985); Komić (2011), of each trigger, relabelling our statistic for $N$ detectors as
@@ V.1 Selecting a time window
As mentioned in Section III.2, we explored different time windows $t_{w}$ for each GW detector, namely H1, L1 and V1. To evaluate the performance of the algorithms without fine-tuning them we use the receiver operator characteristic (ROC) curve, which is represented as the true positive rate (TPR) as a function of false positive rate (FPR) Ting (2010). A relevant point is that our task is a multi-class classification, so to compute the ROC curve, which is usually employed in binary classification, we need to reduce our problem to a pairwise comparison, i.e. Injection class (positive class) against all other classes (negative class).

### v2

@@ II.4 Triggers
The GstLAL pipeline adopted a simple approach to solve this problem: the triggers with the largest SNR peak within the vicinity ($\pm$1\text{\,}\mathrm{s}$\$) will be defined as the cluster centroid and utilized for further analysis [^1] Another clustering of triggers is performed at a later stage of the pipeline. However, this has no importance for our purposes.. While this approach has the benefits of being simple, it discards many pieces of information, such as the position of the various triggers within the template bank and their time ordering. For illustration, Fig. 2 shows the tracks that “matched” GW190521 during the IMBH search in O3 for LIGO Livingston (L1). We represent the IMBH template bank (grey) as a function of the progenitor masses, and the track is coloured as a function of the maximum SNR $\rho$.
Afterwards, each trigger is ranked Cannon:2015gha; Hanna:2019ezx; Tsukada:2023edh according to its probability of originating from an astrophysical signal. The goal of the pipeline is then to obtain a list of triggers, ordered by their likelihood $\Lambda$ to be of astrophysical origin.
@@ III.1 Injections and glitches
Here, we describe the simulated IMBH waveforms, and the six different glitch populations from the data set of Gravity Spy during O3 in H1, L1 and V1 zevin2017gravity; Glanzer:2022avx. As Gravity Spy is a classification algorithm, it returns a classification for each glitch. Glitches have been selected with a Gravity Spy confidence level greater than $90\%$, indicating that the model is highly confident that the glitches belong to the assigned class.
@@
Figure 3: Q-transform of GW190521 and example of different glitch classes selected from Gravity Spy at $90\%$ confidence in LIGO Livingston (L1) zevin2017gravity.
  • Injections: For this aim we use IMRPhenomD approximant Khan:2015jqa with masses $m_{1}\in[50,400]\,\text{M}_{\odot}$ and $m_{2}\in[10,250]\,\text{M}_{\odot}$, and dimensionless spins $\chi_{1},\chi_{2}\in[0,0.99]$, uniformly sampled within the specified range with aligned spins. The distances are uniformly sampled from the range $d\in[10,250]\,$Mpc. The inclination is sampled from a uniform distribution, and the sky-localization is randomized.
  • Blip: these glitches have a characteristic morphology of a symmetric “teardrop” shape in time-frequency in the range [30,250][30,250]\,Hz with short-durations, $\sim 0.04\,$s Cabero:2019orq. Due to their abundance and form, these glitches hinder both the unmodeled burst and modelled CBC can searches with particular emphasis on compact binaries with large total mass due to the short duration of their waveforms, among others Abbott_2018; abbott2016characterization. Moreover, as their source is still unknown, they cannot be removed from astrophysical searches.
  • Fast_\_Scattering: they appear as short-duration arches ($\sim 0.2-0.3\,$s) in the frequency range [2060][20-60]\,Hz. These glitches are strongly correlated with ground motion in range [0.10.3][0.1-0.3]\,Hz and [16][1-6]\,Hz, which in turn is associated with thunderstorms and human activity near the detector. Note that this class was added during O3, and it is more abundant in L1 than H1, but not present in Virgo, due to differences in ground motion and detector sensitivity Soni:2021cjy.
  • Koi_\_Fish: their naming comes from their imaginative frontal resemblance to koi fishes. They are also similar to Blip but typically feature higher-SNR than the latter, spanning the frequency range of $\sim 20-1000\,$Hz.
  • Low_\_Frequency_\_Burst: they are short-duration glitches ($\sim 0.25\,$s) in the frequency range [1020][10-20]\, Hz with a distinctive hump shape at the bottom of spectrgrams. These occurrences were prevalent in L1 data during O1 and H1 data in O3a.
@@
  • Whistle: these glitches have a characteristic V, U or W shape at higher frequencies ($\gtrsim$ 128 Hz) with typical durations $\sim 0.25\,$s. They are caused when radio-frequency signals beat with the voltage-controlled oscillators Nuttall:2015dqa.
@@ III.2 Clustering
    Unknown data set: Usually clusters have a centroid with the highest $\rho$, and a set of neighbours with smaller $\rho$. For illustration, in the top panel of Fig. 4 we can observe the SNR of a cluster with this behavior. In a realistic setting, we do not know the GPS time of the centroid of the cluster $t_{cluster}$ a priori, so we define them following a procedure similar to GstLAL. We order all the triggers according to their GPS time. Afterwards, we divide the whole observing run in windows of $\pm 1\,$s, starting from the beginning of the run. Within each window, we select the trigger with maximum $\rho$ as the centroid, so that $t_{cluster}=t_{\max{\rho}}$. As before, we select $t_{w}=\pm\,\Delta t$ around $t_{cluster}$ to define the cluster.
@@ III.4 Time shift and false alarm rate
$\widetilde{\mathcal{F}}(\Lambda^{*})=\frac{\sum_{l=i}^{L}\widehat{N}_{l}(\Lambda\geq\Lambda^{*})}{T_{b}},$ (7)
@@
$\begin{split}&\widetilde{\mathcal{F}}(\Lambda^{*})=\frac{\sum_{l=i}^{L}\widehat{N}_{l}(\Lambda\geq\Lambda^{*})}{T_{b}T_{f}/T_{f}}\\ &=\frac{\widehat{N}(\Lambda\geq\Lambda^{*})}{T_{f}\widetilde{N}}=\frac{\overline{N}(\Lambda\geq\Lambda^{*})}{T_{f}}.\end{split}$ (8)
@@
Since the offsets of the time shifts should be large enough that each slide can be thought of as an independent experiment, we chose to slide L1 in steps of 33\,s, and V1 in steps of 66\,s with respect to H1, which is fixed. We performed enough time shifts to produce 1,0001,000\,yr of data in triple detector coincidence. Afterwards, the time shifts are binned according to their coincidence type and their time of background.
To minimize the number of false positives, the types of detector time coincidences analyzed in this work are either double or triple. Using Eq. 8, it is necessary to properly measure the time of the search $T_{s}$ and the time of the background $T_{b}$. If we are measuring the time in double-time coincidence, for example, between H1 and L1, it would imply that V1 was down at that time (no V1). Otherwise, if all detectors record data simultaneously, it is considered a triple-time coincidence. In Table 1 we can see an overview of the time of the search $T_{s}$. Since H1L1 (no V1) is extremely small, it is computationally intensive to generate 10001000\,yr of time shifts. Thus, in the interest of time and computational resources we focus on triple-coincident time as we expect a better performance.
@@ IV Methodology
The challenges in GW research require innovative solutions, and ML has emerged as a crucial tool for addressing them due to its adaptability and transversality. The main advantage of ML techniques is their rapidity during the deployment phase, since most of the computations are made during the training stage. In the past few years, researchers have explored different ML applications to GW data analysis, such as detection of CBC George:2016hay; PhysRevD.105.043006; ruder2016overview; Gabbard:2017lja; Sharma:2022ibm; Krastev; Menendez-Vazquez:2020khz; Baltus:2021nme; Baltus:2022pep; 2024arXiv240905068M; Nousi:2022dwh; Koloniari:2024kww, burst identification 2020arXiv200204591C; PhysRevD.105.084054; Iess:2020yqj; Iess:2023quq; LopezPortilla:2020odz; Lopez:2021rci; Meijer:2023yhn; Boudart:2022xib; Boudart:2022apz; Skliris:2020qax, glitch characterization Biswas:2013wfa; zevin2017gravity; Glanzer:2022avx; bahaadini2018machine; Laguarta:2023evo and synthetic data generation Lopez:2022lkd; Lopez:2022dho; Powell:2022pcg; Yan:2022spw; Dooney:2022arh; Dooney:2024pvt, among others. We refer the interested reader to Cuoco:2020ogp and Schafer:2022dxv for a review.
In PhysRevD.105.043006 the authors seek to enhance current state-of-the-art search algorithms with ML applications. Following a similar line of reasoning, and as previously mentioned, we use the triggers generated by the IMBH search of GstLAL during O3 to train a classification algorithm for multiple classes. In this way, we output a statistic to differentiate between IMBH signals and glitches, providing direct information about the nature of these populations.
@@ IV.2 Tackling class imbalance
While we can simulate a large number of GW signals, glitches have a finite nature. Despite their occurrence rate being approximately $\sim 1\,\text{min}^{-1}$, some classes of glitches are more common than others LIGOScientific:2020ibl.
For example, in L1 the Gravity Spy pipeline identified $10^{4}$ Tomtes, but only $\sim 10^{3}$ Whistles. In Table 2 we show the original size of our training and validation set, which is highly imbalanced. This is problematic for classification ML algorithms, as heavily imbalanced datasets tend to be biased towards the majority class to minimize their loss function. This may wrongly be interpreted as a “good” performance, while in reality, the model is only learning a single class: the largest one. Similarly to Powell:2022pcg, to circumvent this issue we use undersampling or oversampling techniques:
@@
    Oversampling: To oversample we use bootstrapping with replacement, which stochastically resamples the existing data set allowing the same example to appear more than once James2013. Using this method, we oversample all glitch classes to equal the size of the largest glitch class.
@@
To avoid overfitting we use K-fold cross-validation journals/jmlr/BengioG04. This method partitions the data set in $K$ subsets or folds (see Fig. 6). The model is evaluated $K$ times, using a different fold as validation in each iteration, while training on the rest of the folds. This allows for a more robust estimate of the model’s performance, as it accounts for the variability in the data and limits the potential for overfitting. In particular, we chose $K=9$ as it is a trade-off between computational complexity and performance.
@@ IV.3 Training model
The unit of an MLP is the perceptron, a mathematical simplification of human neurons rosenblatt1958perceptron. In this unit, the information is given by a $n$-vector $x=[x_{1},\dots,x_{n}]$, that travels in the neuron and interacts multiplicatively with the weights, $w=[w_{1},\dots,w_{n}]$. The goal of the perceptron is to learn the value of the weights, which controls the strength of the influence of the previous neuron. In this simplistic model, the information is summed, such that if the sum is above a certain threshold, the neuron creates or fires a signal. This firing rate is modelled as the activation function $f:\mathbb{R}\rightarrow\mathbb{R}$, which is non-linear. Thus, we can mathematically define the $a$ output, known as activation as,
@@
We implemented the MLP structure in PyTorch NEURIPS2019_9015. After several experiments, we selected the best-performing architecture from our results: it consists of 3 hidden layers of 350 each, using the ReLU activation function relu, except the output layer that uses Softmax activation function, suitable for multi-class classification, since it is a multi-class task. This MLP architecture is common for each GW detector, but they have been trained separately, i.e. it is a single-detector classifier. To optimize the networks we use cross-entropy loss function, and Adam optimizer Kingma:2014vow. Moreover, to adjust the number of epochs and avoid overfitting we implemented an early stopping algorithm Yao2007. We define an epoch as the number of times the network has passed through the whole training and validation data set. Thus, the early stopping algorithm calculates the difference in validation accuracy $\mathcal{A}_{val}$ between the current epoch, $e$, and the best epoch, $e^{\text{best}}$. The training process finishes if
@@
during 150 epochs, with $\epsilon=0.0001$. For the learning rate, we use an adaptive learning rate built-in the PyTorch function ReduceLROnPlateau. Setting the initial learning rate to $10^{-3}$, if the validation accuracy remains constant after 100100\,epochs, the learning rate decreases $10\%$. The combination of these methods proved to increase the performance while decreasing the time needed for fine-tuning.
@@ IV.4 Time coincident tracks
Even though the MLP receives direct information from the Injection class and various glitch classes, misclassifications are still common due to the simplicity of the feature vector (see Section IV.1). Nonetheless, if a potential GW signal is detected in multiple detectors, the likelihood of it being of astronomical origin increases significantly. Hence, it is standard in GW searches to evaluate triggers that happened within the light time travel between detectors: 1010\,ms between H1 and L1, 2727\,ms between H1 and V1, and 2626\,ms between L1 and V1. An additional 55\,ms is considered for statistical fluctuations Sachdev:2019vvd.
@@
In this work, the approach is inherently different, as we do not deal with individual triggers but with averaged tracks. Thus, two tracks from different detectors are considered time coincident if the average time of the triggers within each track falls within the light travel time and time fluctuations. Nevertheless, even if two tracks are time coincident, there is no guarantee that they originate from the same signal. Therefore, two tracks are only considered time coincident if they have at least one common trigger. If tracks from different detectors coincide in time, we assign to the coincidence a probability equal to the harmonic mean, one of Pythagorean means 538b1c7d-3917-388b-be54-a65758913e9f; harmonic, of each trigger, relabelling our statistic for $N$ detectors as
@@ V.1 Selecting a time window
As mentioned in Section III.2, we explored different time windows $t_{w}$ for each GW detector, namely H1, L1 and V1. To evaluate the performance of the algorithms without fine-tuning them we use the receiver operator characteristic (ROC) curve, which is represented as the true positive rate (TPR) as a function of false positive rate (FPR) reference/ml/Ting10. A relevant point is that our task is a multi-class classification, so to compute the ROC curve, which is usually employed in binary classification, we need to reduce our problem to a pairwise comparison, i.e. Injection class (positive class) against all other classes (negative class).